# EgoPlan-Bench2: A Benchmark for Multimodal Large Language Model Planning in Real-World Scenarios

### v1

@@ IV-A Experimental Settings
In this study, we conduct the evaluation on 21 MLLMs, including GPT-4V [3], Video-LLaMA2 [50], ShareGPT4Video [47], LLaVA-NeXT-Video [48], VILA [80], VideoChat2 [18], LongVA [49], Video-LLaVA [79], Video-ChatGPT [40], BLIP-2 [35], InstructBLIP [76], InstructBLIP Vicuna [76], Yi-VL [70], Qwen-VL-Chat [78], Valley [44], DeepSeek-VL [77], LLaVA1.5 [72], mPLUG-Owl-2 [75], MultiModal-GPT [71], InternVL-1.5 [73, 74] and InternVL-2 [73, 74]. For video MLLMs, we adhere to their official configurations, including the number of frames. We crop the task progress video as the visual input and modify the sampling function to ensure inclusion of both the first frame and the last frame (representing the current observation image). For image MLLMs, we consistently use 8 key frames, reducing the number if necessary to prevent inference issues. These key frames are uniformly sampled from the provided video clips and saved in advance for model evaluation.
For the evaluation, we use a common prompt as: Select the best answer to the following multiple-choice question based on the video. Respond with only the letter (A, B, C, or D) of the correct option. Considering the progress shown in the video and my current observation in the last frame, what action should I take next in order to [task goal]? [candidate choices]. Following the evaluation strategy in Video-MME [19], the accuracy is calculated by matching the output of the model with the real one, without introducing any third party model such as GPT.
@@ V Towards Human-level Planning with Multimodal Chain-of-Thought Prompting
In the realm of natural language processing, Chain-of-Thought (CoT) reasoning empowers language models to tackle complex tasks by informing them to generate intermediate rationales. Numerous recent studies [81],  [82],  [83],  [84],  [85] have significantly enhanced the performance of MLLMs through the application of multimodal CoT prompting, which integrates CoT reasoning with additional multimodal prompts.
@@ V-A Prompts for Historical Task Progress
  • Action sequence (Action-seq). Following the long-term memory extraction method described in  [86], we uniformly sample four frames from each action video segment and utilize GPT-4V to condense them into a concise phrase that encapsulates the action content. This process is repeated for all completed actions, resulting in the generation of a text sequence of actions (Action-seq-GPT) that captures the historical task progress. For example, an action sequence might read: “place golf ball on tee, swing the golf club, pick up a golf ball”. We also extract raw annotations to construct a ground-truth action sequence (Action-seq-GT) for comparison purposes.
@@
  • Key objects tracking (Tracking). The motion trajectories of key objects often contain important task information indicating their future movement patterns, which may assist in planning the subsequent action. Inspired by  [83], we initially instruct GPT-4 to identify key objects potentially relevant to the task. An open-vocabulary object detector, Grounding DINO [87], is then used to delineate object bounding boxes in the current observation image. To counter tracking failures due to scene transitions, we crop a four-second video segment and uniformly sample eight frames from it. Additionally, this video segment is reversed to position the current observation image at the beginning of the sequence, focusing the tracking on key objects in the current observation rather than on historical objects. Bounding boxes are utilized as prompts for SAM2  [88] to facilitate video tracking, with consistent coloring employed to identify the same object throughout.
@@ V-B Prompts for Current Observation State
  • Scene graph (SG). Scene graph is a formalization of objects and their relations and attributes that has been extensively used as a bridge between the visual and textual domains. We follow the steps of  [84] to summarize the current observation image into the text-based scene graph.

### v2

@@ IV-A Experimental Settings
In this study, we conduct the evaluation on 21 MLLMs, including GPT-4V [GPT4v], Video-LLaMA2 [cheng2024videollama], ShareGPT4Video [chen2024sharegpt4video], LLaVA-NeXT-Video [zhang2024llavanextvideo], VILA [lin2024vila], VideoChat2 [li2024mvbench], LongVA [zhang2024longva], Video-LLaVA [lin2023video], Video-ChatGPT [maaz2023video], BLIP-2 [li2023blip], InstructBLIP [dai2023instructblip], InstructBLIP Vicuna [dai2023instructblip], Yi-VL [young2024yi], Qwen-VL-Chat [bai2023qwen], Valley [luo2023valley], DeepSeek-VL [lu2024deepseek], LLaVA1.5 [liu2024improved], mPLUG-Owl-2 [ye2024mplug], MultiModal-GPT [gong2023multimodal], InternVL-1.5 [chen2024internvl, chen2024far] and InternVL-2 [chen2024internvl, chen2024far]. For video MLLMs, we adhere to their official configurations, including the number of frames. We crop the task progress video as the visual input and modify the sampling function to ensure inclusion of both the first frame and the last frame (representing the current observation image). For image MLLMs, we consistently use 8 key frames, reducing the number if necessary to prevent inference issues. These key frames are uniformly sampled from the provided video clips and saved in advance for model evaluation.
For the evaluation, we use a common prompt as: Select the best answer to the following multiple-choice question based on the video. Respond with only the letter (A, B, C, or D) of the correct option. Considering the progress shown in the video and my current observation in the last frame, what action should I take next in order to [task goal]? [candidate choices]. Following the evaluation strategy in Video-MME [fu2024video], the accuracy is calculated by matching the output of the model with the real one, without introducing any third party model such as GPT.
@@ V Towards Human-level Planning with Multimodal Chain-of-Thought Prompting
In the realm of natural language processing, Chain-of-Thought (CoT) reasoning empowers language models to tackle complex tasks by informing them to generate intermediate rationales. Numerous recent studies [wang2022language],  [jiang2024joint],  [wang2024vlm],  [mitra2024compositional],  [zhou2024image] have significantly enhanced the performance of MLLMs through the application of multimodal CoT prompting, which integrates CoT reasoning with additional multimodal prompts.
@@ V-A Prompts for Historical Task Progress
  • Action sequence (Action-seq). Following the long-term memory extraction method described in  [shi2024epd], we uniformly sample four frames from each action video segment and utilize GPT-4V to condense them into a concise phrase that encapsulates the action content. This process is repeated for all completed actions, resulting in the generation of a text sequence of actions (Action-seq-GPT) that captures the historical task progress. For example, an action sequence might read: “place golf ball on tee, swing the golf club, pick up a golf ball”. We also extract raw annotations to construct a ground-truth action sequence (Action-seq-GT) for comparison purposes.
@@
  • Key objects tracking (Tracking). The motion trajectories of key objects often contain important task information indicating their future movement patterns, which may assist in planning the subsequent action. Inspired by  [wang2024vlm], we initially instruct GPT-4 to identify key objects potentially relevant to the task. An open-vocabulary object detector, Grounding DINO [liu2023grounding], is then used to delineate object bounding boxes in the current observation image. To counter tracking failures due to scene transitions, we crop a four-second video segment and uniformly sample eight frames from it. Additionally, this video segment is reversed to position the current observation image at the beginning of the sequence, focusing the tracking on key objects in the current observation rather than on historical objects. Bounding boxes are utilized as prompts for SAM2  [ravi2024sam] to facilitate video tracking, with consistent coloring employed to identify the same object throughout.
@@ V-B Prompts for Current Observation State
  • Scene graph (SG). Scene graph is a formalization of objects and their relations and attributes that has been extensively used as a bridge between the visual and textual domains. We follow the steps of  [mitra2024compositional] to summarize the current observation image into the text-based scene graph.
@@ VI Conclusion
[Uncaptioned image] Lu Qiu is a PhD student at Department of Electrical and Electronic Engineering, The University of Hong Kong. She obtained B.Sc. in 2021 and M.Sc. degree in 2024 from Shanghai Jiao Tong University, Shanghai, China. Her research interests include large multimodal models.
[Uncaptioned image] Yuying Ge is a senior researcher at ARC Lab, Tencent PCG. She received her Ph.D. degree from the Multimedia Laboratory (MMLab) at The University of Hong Kong. She has published over 20 papers in top-tier conferences and journals, including CVPR, ICLR, NeurIPS, ECCV, CoRL, ICASSP, and TIP, with a total citation count exceeding 2,000. Her recent representative works include the SEED series and SEED-Bench. Her current research focuses on multimodal representation learning, pretraining, and their downstream applications.
[Uncaptioned image] Yi Chen is a PhD student at Musketeers Foundation Institute of Data Science, The University of Hong Kong. She obtained B.Sc. degree in 2021 and M.Sc. degree in 2024 from Harbin Institute of Technology, Shenzhen, China. Her research interests include large multimodal models and embodied ai.
[Uncaptioned image] Yixiao Ge is a principal researcher at ARC Lab, Tencent PCG, leading on multimodal foundation models, open-world visual comprehension, and efficient AI. She received her Ph.D. degree from the Multimedia Lab (MMLab) at The Chinese University of Hong Kong. She has authored over 50 papers in top-tier conferences and journals, with a total citation count exceeding 5,000. Her recent representative works include SEED, YOLO-World, and LLaMA Pro. She is listed in the 2024 World’s Top 2% Scientists.
[Uncaptioned image] Ying Shan is a Distinguished Scientist at Tencent, the Director of ARC Lab at Tencent PCG, and the Center of Visual Computing at Tencent AI Lab. He has extensive experience in the areas of computer vision, deep learning, and data mining, applying to the full stack of content creation, understanding, and search. Prior to joining Tencent, he worked at Microsoft Bing Ads as a Principal Scientist Manager. He has published more than 180 papers in top conferences and journals and holds many U.S./International patents.
[Uncaptioned image] Xihui Liu (Member, IEEE) is an assistant professor at Department of Electrical and Electronic Engineering and Institute of Data Science, The University of Hong Kong. Before joining HKU, she was a postdoctoral researcher at University of California, Berkeley. She received the Bachelor’s degree from Tsinghua University and PhD degree from The Chinese University of Hong Kong. Her research interests include computer vision, deep learning, generative models, and multimodal AI. She was awarded Adobe Research Fellowship 2020, EECS Rising Stars 2021, and WAIC Rising Star Award 2022. She serves as area chairs for CVPR 2024, ACM MM 2024, ICLR 2025, and CVPR 2025.